### v2

# Constraining Zero-Point Length from Gravitational Baryogenesis

Abstract
The existence of a fundamental zero-point length, $l_{0}$, a minimal spacetime scale predicted by T-duality in string theory or quantum gravity theories, modifies the entropy associated with the horizon of spacetime. In the cosmological setup, this leads to correction to the Friedmann equations governing the evolution of the Universe. In this paper, we investigate the implications of zero-point length $l_{0}$-corrected gravity for gravitational baryogenesis and early universe thermodynamics, deriving constraints on $l_{0}$ from observational baryon asymmetry data. We observe that under the condition of non-equilibrium thermodynamics, $l_{0}$ generates $\dot{\mathcal{R}}\neq 0$ during radiation epoch, where $\mathcal{R}$ is the Ricci scalar. This yields a baryon asymmetry parameter $\eta\propto l_{0}^{2}T_{D}^{9}/M_{\rm Pl}^{7}$. The observed baryon asymmetry $\eta\sim 9.9\times 10^{-11}$ constrains $l_{0}\lesssim 7.1\times 10^{-33}m$, approximately $440$ times the Planck length. Furthermore, our analysis reveals that the zero-point length correction in the Friedmann equation, effectively slows the expansion rate at high energies, resulting in a modified time-temperature relationship where the Universe maintains higher temperatures for longer time during early epochs compared to standard cosmology. Our results establish zero-point length cosmology as a testable framework connecting quantum gravity to cosmological observables, with implications for early universe thermal history and fundamental length scales.
I Introduction
The Friedmann equations, which govern the expansion dynamics of the Universe, are foundational to modern cosmology. These equations emerge from Einstein’s field equations of general relativity (GR) under the assumption of a homogeneous and isotropic universe described by the Friedmann-Robertson-Walker (FRW) metric. However, at extremely high energies approaching the Planck scale, quantum gravitational effects are expected to modify the classical description of spacetime. One such quantum correction arises from the concept of a fundamental zero-point length $l_{0}$, a minimal length scale below which the classical notion of spacetime breaks down. The zero-point length is motivated by various approaches to quantum gravity, including string theory, non-commutative geometry, and generalized uncertainty principles (GUP). For example, in T-duality of string theory, the momentum space propagator for massless particles is given by Padm ; Sma ; Spa ; Fon ; Nicolini
$G(k)=-\frac{l_{0}}{\sqrt{k^{2}}}K_{1}\left(l_{0}\sqrt{k^{2}}\right),$ (1)
where $k=p/\hbar$, $l_{0}$ represents the zero-point length of spacetime, and $K_{1}(x)$ is the modified Bessel function of the second kind. These frameworks suggest that spacetime may possess a discrete or ”fuzzy” structure at microscopic scales, leading to modifications of physical laws near the Planck length ($l_{Pl}\sim 10^{-35}\mathrm{m}$).
Another motivation for considering zero-point length comes from the fact that it is a possible solution to the singularity problem in physics. It is well-known that both GR and Newton’s law of gravitation suffer from the singularity problem. Indeed, the Newtonian gravitation potential for a spherically symmetric mass distribution $M$ is given by $\phi=-GM/r$, which clearly diverges at $r=0$. Based on relation (1), the static Newtonian potential corresponding to $G(k)$ at distance $r$ was shown to take the following form Nicolini ; Nicolini2 ; Nicolini3
$\phi(r)=-\frac{GM}{\sqrt{r^{2}+l_{0}^{2}}},$ (2)
where $M$ is the mass of the source of the potential and for simplicity, we have set $\hbar=c=1$. Thus, a point particle with mass $m$ at distance $r=R$ from mass $M$, feels the force
$\vec{F}=-m\vec{\nabla}\phi(r)|_{r=R}=-\frac{GMm}{R^{2}}\left[1+\frac{l_{0}^{2}}{R^{2}}\right]^{-3/2}\hat{r}$. (3)
Based on the modified Newton’s law of gravitation given in Eq. (3), and applying the notion of the entropic force scenario proposed by Verlinde Ver and developed in Cai2 ; sheyECFE , one can derive the corrections to the entropy associated with the horizon of spacetime KS ; ShLi . These correction terms alter the dynamical field equations of gravity through the thermodynamics-gravity conjecture.
While direct experimental verification of such Planck-scale effects remains elusive, their potential implications for cosmology and early universe physics provide a fertile ground for testing these models. However, cosmological scenarios incorporating a fundamental zero-point length cannot be constrained using observational bounds from Big-Bang nucleosynthesis (BBN), in contrast to analysis such as Refs. Barrow2 ; Anish ; Ava1 ; Ava2 . The reason is that the energy (and corresponding length and time) scales at which zero-point length effects become relevant lie far above the BBN scale, rendering primordial nucleosynthesis completely insensitive to such corrections. Consequently, in this work we employ gravitational baryogenesis which operates at energies many orders of magnitude higher than BBN as a viable probe for constraining zero-point length cosmology Luciv ; Luciv2 ; Luciv3 .
A quintessential cosmological puzzle that may be sensitive to such fundamental physics is the origin of the baryon asymmetry, the observed overwhelming dominance of matter over antimatter in the Universe ref41 . As first outlined by Sakharov ref42 , the generation of a net baryon number requires three necessary conditions: (i) Baryon number (B) violation, (ii) C and CP violation, and (iii) Departure from thermal equilibrium.
The Standard Model of particle physics, while containing sources for all three conditions, is insufficient to generate the observed asymmetry ref41 . Consequently, explanations often lie in extensions to the Standard Model or in new physics at high energies.
Intriguingly, the interplay between gravity and quantum mechanics can provide a natural framework for baryogenesis. A compelling mechanism involves a dynamical breaking of CPT (and consequently CP) symmetry through a coupling between the baryon current $J^{\mu}$ and the derivative of the Ricci scalar curvature $\mathcal{R}$ ref46
$\frac{1}{M_{*}^{2}}\int d^{4}x\sqrt{-g}J^{\mu}\partial_{\mu}\mathcal{R},$ (4)
[…]
This is where the intrinsic structure of spacetime proposed by quantum gravity models becomes crucial. The zero-point length $l_{0}$, implemented via modified horizon entropy and the resulting corrected Friedmann equations, alters the expansion dynamics of the Universe. This modified cosmological evolution implies that the relation between energy density, pressure, and the Hubble parameter is altered, leading to variations in the energy density ($\delta\rho$) and pressure ($\delta p$) from their equilibrium values. These variations signify a deviation from thermal equilibrium, thereby satisfying the third Sakharov condition ref43 ; ref44 ; Luciv .
[…]
This approach connects a fundamental Planck-scale phenomenon directly to an observable cosmological quantity. The measured value of the baryon asymmetry, as constrained by the Cosmic Microwave Background (CMB) anisotropies to be $\eta\sim(6.3\pm 0.3)\times 10^{-10}$ Luciv ; Planck ; ref48 ; ref49 ; ref50 ; ref51 ; ref52 , imposes stringent constraints on the zero-point length parameter $l_{0}$.
In this paper, we explore the implications of the zero-point length model for gravitational baryogenesis. We first derive the modified Friedmann equations by incorporating zero-point length corrections to the entropy of the apparent horizon. We then investigate how this geometric incorporation of a minimal length modifies the cosmological evolution and breaks thermal equilibrium in the early Universe. Our primary goal is to derive the resulting baryon asymmetry and use the observed value of $\eta$ to place observational constraints on the fundamental zero-point length parameter $l_{0}$. By doing so, we aim to bridge the gap between a geometric implementation of a minimal length and a pivotal cosmological observation, offering a potential pathway to test the quantum nature of spacetime.
Our paper is organized as follows. In section II, we review the modified cosmology based on zero-point length correction. In section III, we present the mechanism for gravitational baryogenesis and calculate the resulting baryon asymmetry $\eta$ within our framework and we impose the observational constraint from the measured baryon asymmetry to establish bounds on the zero-point length $l_{0}$. In section IV, we construct the cosmic time as a function of temperature in this cosmological framework and disclose the effects of zero-point on time-temperature relation. The final section is devoted to conclusions.
II Modified cosmology through zero-point length correction
In this section, we review derivation of the modified Friedmann equations inspired by zero-point length corrections to the entropy. We follow ShLi and apply the thermodynamics-gravity conjecture to the apparent horizon of a FRW universe and use the first law of thermodynamics on the apparent horizon. It has been shown that the existence of $l_{0}$ alters the entropy-area relation of black holes as well as the cosmological horizons, which in turn affects the gravitational field equations through the holographic principle. Taking the zero-point length corrections into account, the entropy associated with the apparent horizon is given by ShLi
$S_{h}$ $=$ $\frac{\pi R^{2}}{G}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-1/2}+\frac{3\pi l_{0}^{2}}{G}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-1/2}$ (5)
$-\frac{3\pi l_{0}^{2}}{G}\ln\left(R+\sqrt{R^{2}+l_{0}^{2}}\right).$
While its differential form is ShLi
$dS_{h}=\frac{2\pi R}{G}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-3/2}dR.$ (6)
Although the associated entropy with the apparent horizon is complicated, for $l_{0}^{2}\rightarrow 0$ it recovers the area law of entropy, namely $S_{h}=\pi R^{2}/G$, as expected. It is evident that $l_{0}$ must carry the dimension of length (or inverse energy in natural units), since the corrected entropy-area expression must remain dimensionless in natural units. Note that the area of the apparent horizon is $A=4\pi R^{2}$, while the volume enveloped by the apparent horizon is given by $V=4\pi R^{3}/3$. Note that here $R$ is the apparent horizon radius defined as Hay1 ; Hay2 ; Bak
$R=\frac{1}{\sqrt{H^{2}+k/a^{2}}},$ (7)
where $a(t)$ is the scale factor, $H=\dot{a}/a$ is the Hubble parameter and we assume the background spacetime is given by
$ds^{2}=-dt^{2}+\frac{a^{2}(t)}{1-kr^{2}}dr^{2}+a^{2}(t)r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$ (8)
The temperature associated with the apparent horizon is Hay1 ; Hay2 ; Bak
$T_{h}=\frac{\kappa}{2\pi}=-\frac{1}{2\pi R}\left(1-\frac{\dot{R}}{2HR}\right),$ (9)
where $\kappa$ is the surface gravity. Let us note that for $\dot{R}<2HR$, one may define $T=|\kappa|/2\pi$ to avoid negative temperature. The energy and matter of the universe are in the form of a perfect fluid, $T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu},$ where $\rho$ and $p$ are the energy density and pressure of the matter field inside the universe, respectively. The conservation of energy in the background of FRW cosmology is guaranteed via
$\dot{\rho}+3H(\rho+p)=0.$ (10)
Since the volume of the universe increases due to its eexpansion, thus we have a work term in the first law of thermodynamics. The corresponding work density, in the background of FRW universe filled with a perfect fluid is obtained as Hay2
$W=\frac{1}{2}(\rho-p).$ (11)
Taking the differential of the total energy inside the apparent horizon, $E=\rho(4\pi R^{3}/3)$, we find
$dE=4\pi R^{2}\rho dR-4\pi HR^{3}(\rho+p)dt,$ (12)
where we have used the continuity equation (10). Combining Eqs. (6), (9), (11) and (12) with the first law of thermodynamics on the apparent horizon, $dE=T_{h}dS_{h}+WdV$, after using the continuity equation (10), we arrive at
$\frac{2}{R^{3}}\left(1+\frac{l_{0}^{2}}{R^{2}}\right)^{-3/2}dR=\frac{8\pi G}{3}d\rho.$ (13)
Integrating and then expanding for small values of $l_{0}^{2}/R^{2}$, we reach
$\frac{2}{l_{0}^{2}}\left[1-\frac{l_{0}^{2}}{2R^{2}}+\frac{3}{8}\frac{l_{0}^{4}}{R^{4}}+...\right]+C=\frac{8\pi G}{3}\rho,$ (14)
where $C$ is an integration constant. The standard Friedmann equation is restored for $l_{0}^{2}/R^{2}\rightarrow 0$, provided we choose
$C\equiv\frac{2}{l_{0}^{2}}-\frac{\Lambda}{3},$ (15)
where $\Lambda$ is a constant which can be interpreted as the cosmological constant. Note that one may also define $C=2/l_{0}^{2}$ by absorbing or neglecting the contribution from vacuum energy to the total energy density $\rho$.
Substituting $R$ from Eq. (7), we reach
$H^{2}+\frac{k}{a^{2}}-\alpha\left(H^{2}+\frac{k}{a^{2}}\right)^{2}=\frac{8\pi G}{3}(\rho+\rho_{\Lambda}),$ (16)
where $\alpha=3l_{0}^{2}/4$, $\rho_{\Lambda}=\Lambda/(8\pi G)$. In the absence of the cosmological constant ($\Lambda=0$) and for a flat universe ($k=0$), the modified Friedmann equation is simplified as
$H^{2}-\alpha H^{4}=\frac{8\pi G}{3}\rho.$ (17)
On the other hand, if we define $X=H^{2}+{k}/{a^{2}}$, Eq. (16) can be rewritten as
$\alpha X^{2}+X-\frac{8\pi G}{3}\rho=0,$ (18)
where we have again set $\rho_{\Lambda}=0$, for simplicity. Eq. (18) admits the following solution
$X=\frac{2}{3l_{0}^{2}}\left[1\mp\sqrt{1-8\pi Gl_{0}^{2}\rho}\right].$ (19)
The standard Friedmann equation is recovered when $l_{0}\rightarrow 0$, provided we choose the minus sign in the above equation. Since $l_{0}^{2}$ is very small, we can expand Eq. (19). We find
[…]
The above equation may be rewritten as
$H^{2}+\frac{k}{a^{2}}=\frac{8\pi G}{3}\rho\,(1+\Gamma\rho),$ (21)
where we have defined $\Gamma=2\pi Gl_{0}^{2}$.
Let us note that Eq. (21) in the high energy limit (early universe) where $\Gamma\rho^{2}\gg\rho$ can be written as ($k=0$)
$H^{2}\approx\frac{8\pi G}{3}\Gamma\rho^{2}.$ (22)
Using $p=\rho\omega$ in the continuity equation (10), we reach
$\rho(t)=\rho{{}_{0}}\,a^{-3(1+\omega)}.$ (23)
Since in the early stages, our Universe is dominated by radiation, therefore $\omega=1/3$, and $\rho_{r}=\rho_{r,0}a^{-4}$. In this case Eq. (22) admits a solution of the form $a(t)=C_{1}t^{1/4}$, with $C_{1}\equiv\left[4\sqrt{\frac{8\pi\Gamma}{3}}\rho_{r,0}\right]^{1/4}$. Thus in this case the scale factor evolves as $a(t)\sim t^{1/4}$. Comparing with the radiation dominated era in standard cosmology ($a(t)\sim t^{1/2}$), we see that the rate of the universe expansion is slower when the zero-point length is taken into account in the cosmological field equations.
III Baryogenesis Constraints on Zero-Point Length Cosmology
This section investigates the phenomenological consequences of zero-point length cosmology for baryogenesis, with the objective of confronting the model with observational data and deriving precise constraints on the fundamental scale $l_{0}$.
In this approach, the $l_{0}$-dependent corrections in the modified Friedmann equations are treated as effective modifications to the standard cosmological fluid. We interpret these corrections as generating perturbations in the energy-momentum content of the universe, representing a departure from the perfect thermal equilibrium of standard cosmology ref43 ; Luciv . To quantify these effects, we decompose the total energy density and pressure into their conventional equilibrium components and $l_{0}$-induced perturbations as follows
$\rho$ $=\rho_{0}+\delta\rho\,$, (24)
$p$ $=p_{0}+\delta p\,.$ (25)
We now proceed to quantify the perturbations induced by the zero-point length by substituting the decomposed energy density and pressure, Eqs. (24) and (25), into the modified Friedmann equations. The leading-order corrections are derived as follows.
Beginning with the modified first Friedmann equation
$H^{2}-\frac{3l_{0}^{2}}{4}H^{4}=\frac{8\pi G}{3}\rho.$ (26)
We impose the standard general relativity limit, $H^{2}=\frac{8\pi G}{3}\rho_{0}$, to isolate the $l_{0}$-dependent contributions. This substitution yields
$\frac{8\pi G}{3}\rho_{0}-\frac{3l_{0}^{2}}{4}\left(\frac{8\pi G}{3}\rho_{0}\right)^{2}=\frac{8\pi G}{3}(\rho_{0}+\delta\rho).$ (27)
Solving for the perturbation, we arrive at the leading-order correction to the energy density
$\delta\rho=-2\pi Gl_{0}^{2}\rho_{0}^{2}.$ (28)
A parallel analysis is applied to the second modified Friedmann equation,
$\dot{H}\left(1-\frac{3l_{0}^{2}}{2}H^{2}\right)=-4\pi G(\rho+p).$ (29)
Inserting the decomposed expressions for energy density and pressure from Eq. (24) into the second modified Friedmann equation and employing the previously derived result for the density perturbation $\delta\rho$ in Eq. (28), the corresponding pressure perturbation is determined. This calculation, when performed under the conditions characterizing the radiation-dominated epoch ($\omega=1/3$), yields
$\delta p=-\frac{10\pi Gl_{0}^{2}}{3}\rho_{0}^{2}.$ (30)
Thus, the zero-point length $l_{0}$ manifestly generates non-vanishing perturbations $\delta\rho$ and $\delta p$ in the cosmological fluid. It is evident that in the limit $l_{0}\rightarrow 0$, these perturbations vanish identically, ensuring a smooth recovery of the standard cosmological framework and confirming the consistency of this perturbative treatment.
The perturbations in the energy density and pressure, given by Eqs. (28) and (30), have profound implications for early universe cosmology, particularly for the mechanism of baryogenesis. The observed matter-antimatter asymmetry of the universe, quantified by the baryon-to-photon ratio, stands in stark contrast to the predictions of a universe starting from a perfectly symmetric, thermal state ref41 . The dynamical generation of this asymmetry, known as baryogenesis, is theorized to occur in an expanding and cooling universe. A seminal framework for this process was established by Sakharov ref42 , who identified three necessary criteria: (i) violation of baryon number ($B$) to allow for a net baryon production; (ii) violation of both $C$ and $CP$ symmetries to ensure that the rates for processes producing baryons and antibaryons are different; and (iii) a departure from thermal equilibrium to prevent CPT symmetry from undoing the generated asymmetry.
In the present model, the first two Sakharov conditions are implemented through a well-established gravitational interaction between the baryon current and the spacetime curvature. The crucial third condition departure from thermal equilibrium is naturally satisfied by $l_{0}$-induced perturbations $\delta\rho$ and $\delta p$ ref43 ; ref44 ; Luciv . These perturbations, inherent to the zero-point length corrected cosmology, represent a fundamental departure from the perfect fluid description of the standard radiation-dominated era, thereby providing the necessary out-of-equilibrium environment. This approach aligns with other quantum-gravity-inspired mechanisms, such as those employing the Generalized Uncertainty Principle ref43 or modifications within $f(T)$ gravity ref44 , where the breakdown of standard thermodynamics is also central to generating the baryon asymmetry.
The observed baryon asymmetry of the universe represents one of the most significant outstanding problems in modern cosmology. Gravitational baryogenesis provides a theoretically compelling framework for explaining this matter-antimatter imbalance by connecting it to fundamental spacetime dynamics. This mechanism, which finds its motivation in supergravity theories and effective field theory approaches ref43 ; ref45 , operates through a dynamical breaking of CPT symmetry during cosmic expansion. The fundamental interaction responsible for this effect couples the baryon number current to the gradient of spacetime curvature ref46
$\mathcal{S}_{int}=\frac{1}{M^{2}_{*}}\int d^{4}x\sqrt{-g}J^{\mu}\partial_{\mu}\mathcal{R}.$ (31)
The characteristic energy scale for this interaction is set by $M_{*}=(8\pi G)^{-1/2}\approx 2.4\times 10^{18}$ GeV, corresponding to the reduced Planck mass. This interaction successfully addresses two of the three Sakharov conditions: the derivative coupling to the Ricci scalar ensures the necessary C and CP violation through dynamical CPT breaking, while baryon number violation is assumed to be provided by additional particle physics processes.
In the context of a homogeneous and isotropic universe, this interaction term simplifies significantly, revealing its essential dependence on the temporal evolution of curvature Luciv ; ref46
$\mathcal{S}_{int}=\frac{1}{M^{2}_{*}}(n_{B}-n_{\bar{B}})\dot{\mathcal{R}},$ (32)
where $n_{B}$ and $n_{\bar{B}}$ denote the baryon and antibaryon number densities, respectively. This expression generates an effective chemical potential $\mu_{B}=-\dot{\mathcal{R}}/M^{2}_{*}$ that distinguishes between baryons and antibaryons, with $\mu_{\bar{B}}=-\mu_{B}$. The presence of this potential creates a thermodynamic bias in the early universe’s thermal plasma, preferentially favoring baryonic states over antibaryonic ones. The resulting net baryon number density follows from standard statistical mechanical considerations ref46
$n_{B}-n_{\bar{B}}=\left|\frac{g_{b}}{6}\mu_{B}T^{2}\right|,$ (33)
where $g_{b}\sim\mathcal{O}(1)$ accounts for the internal degrees of freedom of baryonic species. This elegant formulation establishes a direct connection between the time-varying geometry of the universe, characterized by $\dot{R}$, and the generation of baryon asymmetry, providing a geometric origin for one of cosmology’s most fundamental observations.
The baryon asymmetry is conventionally quantified by the observational parameter $\eta$, defined as the net baryon number density normalized by the entropy density ref47
$\frac{\eta}{7}\equiv\frac{n_{B}-n_{\bar{B}}}{s}.$ (34)
This parameter provides a quantitative measure of the net baryon number density relative to the cosmic entropy density, remaining constant throughout the adiabatic expansion of the universe. Substituting the earlier expression for the net baryon density from Eq. (33) and the standard thermodynamic relation for entropy density in the radiation era, $s=\frac{2\pi^{2}}{45}g_{*s}T^{3}$, we derive the fundamental formula for gravitational baryogenesis Luciv ; ref47
$\frac{\eta}{7}=\left|\frac{15g_{b}}{4\pi^{2}g_{*s}}\frac{\dot{\mathcal{R}}}{M^{2}_{*}T}\right|.$ (35)
This crucial expression must be evaluated at the decoupling temperature $T_{D}$, which marks the epoch when baryon number violating interactions freeze out, thereby preserving any generated asymmetry. The parameter $g_{*s}$ represents the effective number of relativistic degrees of freedom contributing to entropy; following established treatments ref47 , we adopt $g_{*s}\approx g_{*}\approx 106$, corresponding to the complete particle content of the Standard Model at high temperatures.
The essential physical insight emerges clearly: a non-vanishing baryon asymmetry requires $\dot{\mathcal{R}}\neq 0$. In conventional cosmology, during radiation domination, the perfect fluid description yields an identically zero Ricci scalar $\mathcal{R}=0$, hence $\dot{\mathcal{R}}=0$, resulting in no baryon asymmetry $\eta=0$. This is precisely where zero-point length cosmology provides the crucial modification: the $l_{0}$-induced perturbations to the energy density and pressure, detailed in Eqs. (24) and (25), create the necessary departure from perfect thermal equilibrium required by Sakharov’s third condition.
[…]
$\mathcal{R}=-8\pi GT_{g},$ (36)
where $T_{g}=\rho-3p$ is the trace of the energy-momentum tensor. Incorporating the $l_{0}$-corrected density and pressure perturbations from Eqs. (28) and (30), the trace becomes
$T_{g}$ $=(\rho_{0}+\delta\rho)-3(p_{0}+\delta p)$
$=\rho_{0}-3p_{0}+\delta\rho-3\delta p$
$=\rho_{0}-3p_{0}-2\pi Gl_{0}^{2}\rho_{0}^{2}+10\pi Gl_{0}^{2}\rho_{0}^{2}$
$=T_{0}+8\pi Gl_{0}^{2}\rho_{0}^{2}.$ (37)
Substituting this non-zero trace into the field equation yields the modified Ricci scalar
$\mathcal{\mathcal{R}}_{z}=\mathcal{R}_{0}-64\pi^{2}G^{2}l_{0}^{2}\rho_{0}^{2},$ (38)
where $\mathcal{\mathcal{R}}_{0}=-8\pi G(\rho_{0}-3p_{0})$ is the GR Ricci scalar. For the radiation epoch ($\omega=1/3$) we have $\mathcal{R}_{0}=0$ as it is expected.
Taking the time derivative and employing the continuity equation for radiation $\dot{\rho}_{0}=-4H\rho_{0}$ along with the standard Friedmann relation $H^{2}=\frac{8\pi G}{3}\rho_{0}$, we derive the pivotal result
$\dot{\mathcal{R}}_{z}=512\pi^{2}G^{2}l_{0}^{2}H\rho_{0}^{2}\neq 0.$ (39)
The derived expression for the time evolution of the Ricci scalar $\dot{\mathcal{R}}_{z}\propto l_{0}^{2}$ , demonstrates that a nonzero zero-point length directly drives curvature dynamics in the early universe, with the standard cosmological result $\dot{\mathcal{R}}=0$ emerging smoothly in the limit $l_{0}\to 0$. Consequently, the magnitude of the generated baryon asymmetry, being linearly proportional to $\dot{\mathcal{R}}_{z}$, increases with $l_{0}$. This direct correlation identifies $l_{0}$ as the key parameter governing the departure from thermal equilibrium and the resulting matter-antimatter asymmetry. We now substitute Eq. (39) into the expression for $\eta$. Using the energy density of relativistic particles $\rho_{0}={\pi^{2}g_{*}T^{4}}/{30}$ and the Planck mass $M_{*}^{2}=({8\pi G}^{-1})$, a detailed calculation yields the fundamental scaling relation
$\eta=\frac{7\pi^{3}}{\sqrt{3}\cdot 30^{3/2}}\cdot g_{*}^{3/2}\cdot\frac{l_{0}^{2}T_{D}^{9}}{M_{\rm Pl}^{7}}.$ (40)
The numerical evaluation gives the more transparent expression
$\eta\approx 0.762\cdot g_{*}^{3/2}\cdot\frac{l_{0}^{2}T_{D}^{9}}{M_{\rm Pl}^{7}},$ (41)
where we have set the decoupling temperature $T_{D}=M_{I}\approx 3.3\times 10^{16}$ GeV, consistent with upper bounds from inflationary tensor modes ref43 ; ref46 .
Confronting this prediction with the observed value $\eta_{obs}\lesssim 9.9\times 10^{-11}$ ref48 ; ref49 ; ref50 ; ref51 ; ref52 , we obtain a stringent constraint on the zero-point length (see Fig. 1)
$l_{0}\lesssim 3.576\times 10^{-17}\ \text{GeV}^{-1}\approx 7.1\times 10^{-33}\ \text{m}.$ (42)
Expressed in natural units relative to the Planck length $l_{P}=1.616\times 10^{-35}$ m
$\frac{l_{0}}{l_{P}}\lesssim 440.$ (43)
Figure 1: The predicted baryon asymmetry $\eta$ as a function of the zero-point length $l_{0}$. The horizontal dashed line marks the observational upper limit. The viable parameter space, consistent with measurements, is the area beneath this line ($\eta<9.9\times 10^{-11}$).
This bound not only validates our perturbative treatment ($l_{0}\ll 1$) but also suggests that any fundamental minimal length in nature must be proximate to the Planck scale. The characteristic scaling $\eta\propto l_{0}^{2}T_{D}^{9}/M_{\rm Pl}^{7}$ reveals the profound sensitivity of gravitational baryogenesis to high-energy physics and establishes a concrete connection between quantum gravitational effects and cosmological observables.
IV The relation between cosmic time and temperature in zero-point length cosmology
This section analyzes how the zero-point length cosmology modifies the fundamental connection between cosmic time and temperature in the early universe. The zero-point-corrected Friedmann equations alter the high-energy expansion dynamics and thereby lead to a nonstandard thermodynamic evolution. As a consequence, the time-temperature relation deviates from its familiar general-relativistic form.
During the radiation-dominated phase, the primordial plasma remains in thermal equilibrium, implying conservation of entropy within any comoving volume. This standard thermodynamic requirement leads to the condition Weinberg
$s(T)a^{3}=\text{const},$ (44)
where $s(T)$ represents the entropy density. The time derivative of Eq. (44) yields the differential equation governing entropy evolution
$\dot{s}(T)a^{3}+3\dot{a}a^{2}s(T)=0\quad\Rightarrow\quad\frac{ds(T)}{dt}a^{3}=-3\dot{a}a^{2}s(T).$ (45)
Inserting the modified Hubble parameter in SI units
$H=H_{\rm GR}\left(1+\frac{\pi Gl_{0}^{2}}{c^{4}}\rho\right),$ (46)
into Eq. (45) leads to
$\frac{ds(T)}{dt}=-3\left[H_{\text{GR}}\left(1+\frac{\pi Gl_{0}^{2}}{c^{4}}\rho\right)\right]s(T).$ (47)
This equation can be rearranged to express the differential time element as
$dt=-\frac{ds(T)}{3s(T)}\left[H_{\text{GR}}\left(1+\frac{\pi Gl_{0}^{2}}{c^{4}}\rho\right)\right]^{-1}.$ (48)
The resulting functional connection between cosmic time $t$ and temperature $T$ is derived through integration of the above expression
$t=-\frac{1}{3}\int\frac{s^{\prime}(T)}{s(T)}\left[H_{\text{GR}}\left(1+\frac{\pi Gl_{0}^{2}}{c^{4}}\rho\right)\right]^{-1}dT,$ (49)
where the prime notation indicates differentiation with respect to temperature $T$. In the radiation-dominated regime, characterized by $p=\rho/3$, the entropy density and energy density follow the standard thermodynamic expressions Weinberg
$s(T)$ $=\frac{2\mathcal{N}a_{B}T^{3}}{3},$ (50)
$\rho(T)$ $=\frac{\mathcal{N}a_{B}T^{4}}{2},$ (51)
where $\mathcal{N}$ enumerates the total number of relativistic degrees of freedom, accounting for all particles and antiparticles with distinct spin states counted separately Weinberg . Using $\frac{s^{\prime}(T)}{s(T)}=\frac{3}{T}$ in Eq. (49), and employing a Taylor expansion, we obtain
$t=-\int\frac{1}{TH_{\text{GR}}}\left(1-\frac{\pi Gl_{0}^{2}}{c^{4}}\rho\right)dT.$ (52)
Substituting the expressions for $\rho(T)$ and $H_{\text{GR}}=\sqrt{{8\pi G}/{3c^{2}}\rho}$ into Eq. (52) yields
$t=-\int\left(\frac{1}{\beta T^{3}}-\frac{\pi G\mathcal{N}a_{B}l_{0}^{2}}{2c^{4}\beta}T\right)dT,$ (53)
where we have defined $\beta\equiv\sqrt{\frac{8\pi G\mathcal{N}a_{B}}{6c^{2}}}$. Performing the integration, we obtain
$t=\frac{1}{2\beta T^{2}}+\frac{\pi G\mathcal{N}a_{B}}{4c^{4}\beta}l_{0}^{2}T^{2}+\text{const}.$ (54)
This result can be reformulated as
$t=\frac{1}{2\beta T^{2}}\left(1+\frac{3\beta^{2}}{8c^{2}}l_{0}^{2}T^{4}\right)+\text{const}.$ (55)
Expressing this result in terms of the zero-point length parameter $\alpha=\frac{3l_{0}^{2}}{4c^{2}}$ and in SI units, we arrive at the form
$t=\frac{1}{2\beta T^{2}}\left(1+\frac{\alpha}{2}\beta^{2}T^{4}\right)+\text{const},$ (56)
which is the time-temperature relation in zero-point length cosmology. In the limiting case where $l_{0}\to 0$ ($\alpha=0$), the time-temperature relation simplifies to $t={1}/{2\beta T^{2}}$, which corresponds exactly to the standard general relativity result Weinberg . We note that while our modified Friedmann equations were originally derived in natural units, dimensional consistency necessitates that the right-hand side of Eq. (56) must possess dimensions of time $[t]$, and accordingly all fundamental constants have been properly restored.
For the early universe’s hot plasma composition, which includes photons, three generations of neutrinos, antineutrinos, and electron-positron pairs, the total effective relativistic degrees of freedom amount to $\mathcal{N}=43/4$. Consequently, in centimeter-gram-second (cgs) units, Eq. (56) becomes
$t=0.994\left(\frac{T}{10^{10}K}\right)^{-2}F(l_{0},T)+\text{const},$ (57)
where the modification factor $F(l_{0},T)$ is defined as
$F(l_{0},T)\equiv 1+4.46\times 10^{-39}\ l_{0}^{2}T^{4}.$ (58)
The $l_{0}\to 0$ limit eliminates all modifications ($F\to 1$), guaranteeing consistency with standard GR Weinberg . Fig. 2 displays the evolution of temperature $T$ versus cosmic time $t$ across various values of the re-scaled zero-point length parameter $\tilde{l}_{0}\equiv 10^{24}\;l_{0}$, demonstrating that larger values of zero-point length cause higher temperatures in the early universe.
Figure 2: The behavior of the temperature vs time in the radiation dominated era of zero-point length cosmology for different values of the zero-point length parameter $\tilde{l}_{0}=[5\times 10^{-10},7\times 10^{-10},10^{-9}]$.
This thermal enhancement arises from the modified expansion dynamics: the zero-point length correction term ${3l_{0}^{2}}H^{4}/4$ in the Friedmann equation effectively slows the expansion rate at very high energies, allowing the universe to maintain higher temperatures for longer durations compared to standard cosmology. These results highlight the distinctive thermodynamic signatures of zero-point length cosmology in the early universe.
V Conclusions
The introduction of a fundamental zero-point length $l_{0}$ as a signature of quantum gravity leads to measurable modifications in baryogenesis predictions. By incorporating the $l_{0}$-corrected entropy expression associated with the apparent horizon and applying the first law of thermodynamics, we derived the modified Friedmann equations that govern the cosmic expansion history. These geometrically induced corrections were essential for establishing the altered time-temperature relation in the early universe.
[…]
$\eta\propto l_{0}^{\,2}T_{D}^{9}/M_{\rm Pl}^{7}.$
By comparing this theoretical prediction with the observational value $\eta_{\rm obs}\simeq 9.9\times 10^{-11}$, we derived a stringent upper limit on the zero-point length,
$l_{0}\lesssim 7.1\times 10^{-33}\,\text{m},$
which corresponds to roughly $440$ times the Planck length.
We further analyzed the influence of the $l_{0}$-dependent term in the modified Friedmann equation on the early-universe thermal evolution and derived the corresponding time-temperature relation for zero-point length cosmology. Our analysis shows that the zero-point length alters the cosmic cooling rate, yielding $t={1}/{(2\beta T^{2})}\left(1+{\alpha}\beta^{2}T^{4}/2\right).$ This expression demonstrates that, for any fixed cosmic time, our Universe retains a higher temperature than predicted by standard GR. Physically, this behavior arises because the correction term ${3l_{0}^{2}}H^{4}/4$ in the modified Friedmann equation slows the expansion rate at high energies, thereby reducing the cooling rate and allowing the universe to remain hotter for an extended duration. Such a modified thermal history affects several early-universe processes, including phase transitions and relic-abundance calculations.
The consistency between the constraint we found on $l_{0}$ and expectations from quantum gravity, together with the modified thermal history implied by the $l_{0}$-corrected Friedmann dynamics, suggests that zero-point length cosmology provides a coherent and testable framework for probing quantum-gravitational effects. The baryon asymmetry constraint derived here constitutes a measurable imprint of such microscopic physics and indicate that any fundamental minimal length must be extremely small yet potentially accessible to future precision cosmological observations.
It is important to note that, in addition to the zero-point length correction to the area law of entropy discussed in this work, various generalization of the Bekenstein-Hawking entropy has been explored in the literatures Barrow ; Tsallis ; Kan1 ; Kan2 . While the cosmological consequences of these modified entropy (Barrow ; Tsallis ; Kan1 ; Kan2 ) on the BBN at the early stages of the Universe have been explored in Barrow2 ; Anish ; Ava1 ; Ava2 , the influences of the Kaniadakis, Tsallis and Barrow entropies on the early baryogenesis have been studied in Luciv ; Luciv2 ; Luciv3 . Therefore, it is worth constraining other modified entropies through gravitational baryogenesis. These issues are now under investigations and the results will be appeared elsewhere.
Acknowledgements.
We are grateful to Shiraz university Research Council. The work of A. Sheykhi is based upon research funded by Iran National Science Foundation (INSF) under project No. 4022705.